OpenStudy (anonymous):

Find the exact value of the expression. 8sec(pi/6) + 7cot(pi/3)

OpenStudy (anonymous):

= 16/sqrt3 + 7sqrt3/3

OpenStudy (mertsj):

yes That is correct

OpenStudy (anonymous):

This is right, but you can simplify a little more. You can reduce this to \[\frac{16}{\sqrt{3}}+\frac{7\sqrt{3}}{3}=\frac{16\sqrt{3}}{3}+\frac{7\sqrt{3}}{3}=\frac{23\sqrt{3}}{3}\]

OpenStudy (mertsj):

But you could simplify it further.

OpenStudy (anonymous):

Excellent, thank you.